OpenStudy (anonymous):

complete the square and then identify the ordered pair (h.k) 2y=5-3x^2-3x

OpenStudy (anonymous):

is (h,k) the vertex?

OpenStudy (anonymous):

you are going to have to write this as \[y=\text{expression}\] as a first step, so divide by 2

OpenStudy (anonymous):

the instructions for the problem are : for the following equation, complete the square and express the answer in appropriate form, either y=a(x-h)^2+k or x=a(y-k)^2+h. Then identify the ordered pair (h.k). i know how to complete squares, but then?

OpenStudy (anonymous):

\[y=-\frac{3}{2}x^2-\frac{3}{2}x+\frac{5}{2}\]

OpenStudy (anonymous):

annoying but there is no way around it

OpenStudy (anonymous):

ok im following

OpenStudy (anonymous):

then you job is to rewrite as \[y=a(x-h)^2+k\] and pretty clearly \[a=-\frac{3}{2}\]

OpenStudy (anonymous):

question is what is h, and you have a choice on how to compute it you could write \[y=-\frac{3}{2}(x^2+x)+\frac{5}{2}\] as a first step and then write \[y=-\frac{3}{2}(x+\frac{1}{2})^2+\text{stuff}\] i will explain how to find the "stuff" easily

OpenStudy (anonymous):

oh no please no myininaya

OpenStudy (anonymous):

or you could do it the easy way and say the vertex is at \[x=-\frac{b}{2a}=-\frac{1}{2}\] and be done

myininaya (myininaya):

\[y=-\frac{3}{2}(x^2+x)+\frac{5}{2}=\frac{-3}{2}(x^2+x+(\frac{1}{2})^2)+\frac{5}{2}+\frac{3}{2}\cdot (\frac{1}{2})^2\] \[=\frac{-3}{2}(x+\frac{1}{2})^2+\frac{5}{2}+\frac{3}{8}\]

myininaya (myininaya):

\[=\frac{-3}{2}(x+\frac{1}{2})^2+\frac{23}{8}\]

OpenStudy (anonymous):

now that i know the vertex has x coordinate \[-\frac{1}{2}\] i know it looks like \[y=-\frac{3}{2}(x+\frac{1}{2})^2+k\] and k is what i get when i replace x by \[\frac{1}{2}\] in the original equation. you will get \[\frac{23}{8}\]

OpenStudy (anonymous):

if you can follow all that adding and subtracting an keeping track by my esteemed colleague myininaya that way will work. i much prefer computing \[-\frac{b}{2a}\] and then the vertex is \[(-\frac{b}{2a},k)\] where k is what you get when you replace x by \[-\frac{b}{2a}\] in the original expression

OpenStudy (anonymous):

your choice

OpenStudy (anonymous):

i know which one you choose!

myininaya (myininaya):

\[k=f(-\frac{b}{2a})\]

OpenStudy (anonymous):

how was dinner?

myininaya (myininaya):

yummy :)

OpenStudy (anonymous):

oooh yes! that is what i use.

myininaya (myininaya):

but i guess you could just where k is what you get when you replace x by -b/(2a) but thats so long when you can just say f(-b/(2a))

myininaya (myininaya):

:)